# Dielectronic recombination studies of ions relevant to kilonovae and non-LTE plasma

### v1

@@ III Results and Discussion
Figure 5 illustrates the dependence of the total DR rate coefficients on electron temperature, calculated by incorporating Rydberg states with principal quantum numbers ranging from $n=6$ to $n=35$, except for $\text{Ce}^{2+}$. While the high-temperature rates ($T>10^{5}$ K) may not have significant practical relevance, they are included for completeness. At high temperatures $T$, the DR rate coefficient decreases following a $\sim T^{-3/2}$ dependence, as the increasing kinetic energy of free electrons reduces the probability of recombination. In contrast, at low temperatures, the DR process is predominantly influenced by a number of resonances situated just above the threshold. These resonances contribute significantly to the recombination rate, ensuring that DR remains efficient even at lower temperatures. This occurs because the $\sim T^{-3/2}$ factor compensates for the lowness of the temperature. The calculated values have an estimated uncertainty of at least 38%, determined using the error propagation rule from Eq. II. At low temperatures, DR is primarily governed by near-zero-energy resonances, where the energy of free electrons closely matches the energy levels of states within the ion, enabling highly efficient recombination processes. The dominance of these resonances enhances the recombination probability, significantly influencing the ionization balance of the kilonova ejecta. Notably, the behavior of DR at these temperatures can resemble radiative recombination, highlighting the need for precise resonance data for accurate modeling. As the temperature increases, contributions from high-$n$ Rydberg states become more prominent. These states, with their highly excited electrons near the ionization threshold, play a crucial role in recombination dynamics at higher temperatures. The interplay between the thermal energy of the electrons and the populations of these excited states alters the overall recombination efficiency, resulting in a shift in the effective ionization balance. This shift is particularly significant for ions with one or two electrons outside closed shells (e.g., $\text{Y}^{+}$, $\text{Sr}^{+}$, and $\text{Ce}^{2+}$), as it impacts the interpretation of kilonova spectra and the modeling of non-LTE plasmas [7].
Table 1: DR rate-coefficients at temperature $10^{4}$ K.
Ion Rate-coefficient ($10^{-11}\,\text{cm}^{3}/\text{s}$)
Y+ 9.61403
Sr+ 0.21363
Te2+ 0.31646
Ce2+ 10.96778
@@
Among the studied ions, $\text{Ce}^{2+}$ exhibits relatively high DR rate coefficients at very low temperatures compared to the other ions. In the kilonova-relevant temperature range ($\sim 10^{4}$ K), as shown in Table 1, the rate coefficients for $\text{Y}^{+}$ and $\text{Ce}^{2+}$ become comparable and significantly exceed those of $\text{Sr}^{+}$ and $\text{Te}^{2+}$. This indicates that $\text{Y}^{+}$ and $\text{Ce}^{2+}$ have a stronger tendency to populate lower ionization states over time, while $\text{Sr}^{+}$ and $\text{Te}^{2+}$ remain ionized for a longer duration in the expanding ejecta. These DR trends have critical implications for kilonova spectra. Higher DR rates, as seen for $\text{Y}^{+}$ and $\text{Ce}^{2+}$, are expected to promote efficient recombination to lower ionization states, contributing to optical and near-infrared spectral features. In contrast, the slower recombination associated with $\text{Sr}^{+}$ and $\text{Te}^{2+}$ can lead to the retention of higher ionization states, producing spectral features characterized by higher-energy transitions and shifts toward shorter wavelengths. The availability of these computed DR rates will be instrumental in refining spectral models, improving predictions of line emission, and enhancing our understanding of kilonova ejecta composition, as they provide crucial inputs for astrophysical modeling tools like the SUMO spectral synthesis code used in kilonovae simulations [47].
@@
In Figure 6, a comparison is made between the DR rate coefficients obtained in this work and those reported in the only existing study over the temperature range characteristic of the nebular phase of kilonovae. During this phase, the temperature is typically around 10,000 K [48, 49]. To account for the temperature evolution from the early weeks after the event to later stages, influenced by the evolving ejecta composition, the rate coefficients are presented across a broader temperature range from 1,000 K to 30,000 K.
In the case of $\text{Y}^{+}$, the present values are approximately an order of magnitude larger than those reported by Banerjee et al. [36]. The same number of intermediate states has been considered in both studies; however, the difference in magnitude arises from the inclusion of a significantly larger number of Rydberg states in the present work. The inclusion of high-$n$ Rydberg states enhances the recombination probability [50]. Additionally, differences in the number of orbital angular momentum states ($l$) considered in the two studies may further contribute to the observed discrepancy in the rate coefficients.
A different trend is observed for $\text{Sr}^{+}$, where the present results show good agreement with those of Banerjee et al. [36]. This consistency can be attributed to the inclusion of an additional $4p^{6}4f$ state in their study, which increases their rate coefficient. However, this enhancement is counterbalanced by their lower number of Rydberg states and $l$-values, resulting in an overall accidental numerical agreement between the two datasets. The role of Rydberg states is particularly significant in singly charged ions like $\text{Y}^{+}$ and $\text{Sr}^{+}$, where dense energy levels allow for extensive recombination pathways [51]. Despite some variations in magnitude, the present DR rate coefficients remain within an order of magnitude—or better—compared to previous studies. This level of consistency supports the reliability of the present work.

### v2

@@ III Results and discussion
Figure 5 illustrates the dependence of the total DR rate coefficients on electron temperature, calculated by incorporating Rydberg states with principal quantum numbers ranging from $n=6$ to $n=35$, except for $\text{Ce}^{2+}$. While the high-temperature rates ($T>10^{5}$ K) may not have significant practical relevance, they are included for completeness. At high temperatures $T$, the DR rate coefficient decreases following a $\sim T^{-3/2}$ dependence, as the increasing kinetic energy of free electrons reduces the probability of recombination. In contrast, at low temperatures, the DR process is predominantly influenced by a number of resonances situated just above the threshold. These resonances contribute significantly to the recombination rate, ensuring that DR remains efficient even at lower temperatures. This occurs because the $\sim T^{-3/2}$ factor compensates for the lowness of the temperature. The calculated values have an estimated uncertainty of around 38%, determined using the error propagation rule from Eq. II. At low temperatures, DR is primarily governed by near-zero-energy resonances, where the energy of free electrons closely matches the energy levels of states within the ion, enabling highly efficient recombination processes. The dominance of these resonances enhances the recombination probability, significantly influencing the ionization balance of the kilonova ejecta. Notably, the behavior of DR at these temperatures can resemble radiative recombination, highlighting the need for precise resonance data for accurate modeling. As the temperature increases, contributions from high-$n$ Rydberg states become more prominent. These states, with their highly excited electrons near the ionization threshold, play a crucial role in recombination dynamics at higher temperatures. The interplay between the thermal energy of the electrons and the populations of these excited states alters the overall recombination efficiency, resulting in a shift in the effective ionization balance. This shift is particularly significant for ions with one or two electrons outside closed shells (e.g., $\text{Y}^{+}$, $\text{Sr}^{+}$, and $\text{Ce}^{2+}$), as it impacts the interpretation of kilonova spectra and the modeling of non-LTE plasmas (7).
It is worth noting that the uncertainty reported in the present work may appear large; however, we have included an additional 10% uncertainty to account for the contribution from higher Rydberg states that were not included in the present calculations. If experimental observations are carried out for the current set of ions, it is highly likely that these contributions would also be suppressed due to field ionization effects, which typically limit the population of high-$n$ Rydberg states in experimental setups. Therefore, the additional 10% uncertainty we have included would have a relatively minor impact when comparing with experimental results. Nonetheless, we have included this 10% additional uncertainty to provide a more cautious upper bound on the theoretical uncertainty. Furthermore, several previous studies on both high- and low-charge ions (47, 48, 45) have reported similarly large uncertainties, both in terms of energy scale and rate coefficient, which is to be expected when dealing with complex, many-electron systems such as those considered in this work.
Table 2: DR rate coefficients ($10^{-11}\,\text{cm}^{3}/\text{s}$) at temperatures $10^{3}$, $10^{4}$, and $10^{5}$ K.
Ion $10^{3}$ K $10^{4}$ K $10^{5}$ K
Y+ 17.02138 9.61403 1.08042
Sr+ 0.13610 0.21363 0.07922
Te2+ 2.48856 0.31646 1.52769
Ce2+ 91.13522 10.96778 0.48809
@@
Table 2 presents the DR rate coefficients at temperatures of $10^{3}$, $10^{4}$, and $10^{5}$ K. The complete set of DR rate coefficients as a function of temperature is provided in the supplementary material. Among the studied ions, $\text{Ce}^{2+}$ exhibits relatively high DR rate coefficients at very low temperatures compared to the other ions. In the kilonova-relevant temperature range ($\sim 10^{4}$ K), as shown in Table 2, the rate coefficients for $\text{Y}^{+}$ and $\text{Ce}^{2+}$ become comparable and significantly exceed those of $\text{Sr}^{+}$ and $\text{Te}^{2+}$. This indicates that $\text{Y}^{+}$ and $\text{Ce}^{2+}$ have a stronger tendency to populate lower ionization states over time, while $\text{Sr}^{+}$ and $\text{Te}^{2+}$ remain ionized for a longer duration in the expanding ejecta. These DR trends have critical implications for kilonova spectra. Higher DR rates, as seen for $\text{Y}^{+}$ and $\text{Ce}^{2+}$, are expected to promote efficient recombination to lower ionization states, contributing to optical and near-infrared spectral features. In contrast, the slower recombination associated with $\text{Sr}^{+}$ and $\text{Te}^{2+}$ can lead to the retention of higher ionization states, producing spectral features characterized by higher-energy transitions and shifts toward shorter wavelengths. The availability of these computed DR rates will be instrumental in refining spectral models, improving predictions of line emission, and enhancing our understanding of kilonova ejecta composition, as they provide crucial inputs for astrophysical modeling tools like the SUMO spectral synthesis code used in kilonovae simulations (49).
@@
In Figure 6, a comparison is made between the DR rate coefficients obtained in this work and those reported in the only existing study over the temperature range characteristic of the nebular phase of kilonovae. During this phase, the temperature is typically around 10,000 K (50, 51). To account for the temperature evolution from the early weeks after the event to later stages, influenced by the evolving ejecta composition, the rate coefficients are presented across a broader temperature range from 1,000 K to 30,000 K.
In the case of $\text{Y}^{+}$, the present values are approximately an order of magnitude larger than those reported by Banerjee et al. [33]. The same number of intermediate states has been considered in both studies; however, the difference in magnitude arises from the inclusion of a significantly larger number of Rydberg states in the present work. The inclusion of high-$n$ Rydberg states enhances the recombination probability (52). Additionally, differences in the number of orbital angular momentum states ($l$) considered in the two studies may further contribute to the observed discrepancy in the rate coefficients.
A different trend is observed for $\text{Sr}^{+}$, where the present results show good agreement with those of Banerjee et al. [33]. This consistency can be attributed to the inclusion of an additional $4p^{6}4f$ state in their study, which increases their rate coefficient. However, this enhancement is counterbalanced by their lower number of Rydberg states and $l$-values, resulting in an overall accidental numerical agreement between the two datasets. The role of Rydberg states is particularly significant in singly charged ions like $\text{Y}^{+}$ and $\text{Sr}^{+}$, where dense energy levels allow for extensive recombination pathways (53). Despite some variations in magnitude, the present DR rate coefficients remain within an order of magnitude—or better—compared to previous studies. This level of consistency supports the reliability of the present work.
@@ IV Conclusions
V Data availability
The supplementary data set is available at DOI: 10.5281/zenodo.15683884.